# A causal random synthetic turbulent velocity field: I. The underlying Gaussian structure

### v1

@@ 3.2.1 Fourier modes of a white noise
and
$\mathbb{E}\left[\widehat{{\mathrm{d}}W}_{i}(t,\mathbf{k}_{\mathbf{n}})\overline{\widehat{{\mathrm{d}}W}_{j}(t^{\prime},\mathbf{k}_{\mathbf{m}})}\right]=L^{3}_{\text{\tiny{tot}}}\delta_{i,j}\delta_{\mathbf{n},\mathbf{m}}^{(3)}\delta(t-t^{\prime})\ {\mathrm{d}}t\ {\mathrm{d}}t^{\prime},$ (71)
@@ 3.2.2 The statistical structure of the stationary solution: spatial Fractional Gaussian vector Field
Considering the spatial white noise ${\mathrm{d}}W(\mathbf{x})$”, with statistical structure provided in (64) and (3.2.1), and its Fourier modes $\widehat{{\mathrm{d}}W}(\mathbf{k}_{\mathbf{n}})$ with covariance (67) and (68), the real random field $\mathbf{u}^{\text{\tiny{fGv}},\nu}(\mathbf{x})$ is fully defined by its Fourier mode, and is given, component-wise, by
@@
for a given discrete wave vector $\mathbf{k}=\mathbf{k}_{\mathbf{n}}$. $T_{\mathbf{k}}$ is the characteristic time of the Fourier mode $\mathbf{k}$ defined in (56), $\widehat{P}_{ij}(\mathbf{k})$ Leray projector ensuring the incompressibility defined in (7) and $E_{\nu}^{\text{\tiny{E}}}(k)$ is the previously defined 3D-PSD obtained from (54) using the isotropic relation (9). The superscript (1) appearing in the notation of the resulting velocity field $\widehat{u}^{(1),\nu}_{i}(t,\mathbf{k})$ (74) illustrates that it is obtained using a single evolution with respect to the white noise. The justification for this notation will become clear in the next paragraph. The random contribution to the evolution (74) is provided by the stochastic vector term $\widehat{{\mathrm{d}}W}_{p}(t,\mathbf{k})$ characterized by the covariances (70) and (71).
The unique solution of the stochastic dynamics proposed in (74) reads
$\widehat{u}^{(1),\nu}_{i}(t,\mathbf{k})=\widehat{u}^{(1),\nu}_{i}(0,\mathbf{k})e^{-\frac{t}{T_{\mathbf{k}}}}+\sqrt{\frac{E_{\nu}^{\text{\tiny{E}}}(\mathbf{k})}{T_{\mathbf{k}}}}\widehat{P}_{ip}(\mathbf{k})\int_{s=0}^{t}e^{-\frac{t-s}{T_{\mathbf{k}}}}\widehat{{\mathrm{d}}W}_{p}(s,\mathbf{k}),$ (75)
for a given initial condition $\widehat{u}^{(1),\nu}_{i}(0,\mathbf{k})$. It is well known that the Ornstein-Uhlenbeck dynamics proposed in (74) eventually reaches a statistically stationary regime at large time. Moreover, the solution $\widehat{u}^{\nu}_{i}(t,\mathbf{k})$ will get independent of the initial condition at large time, as it can be seen from its expression provided in (75) where the initial condition $\widehat{u}^{(1),\nu}_{i}(0,\mathbf{k})$ is weighted by an exponentially decreasing factor.
It turns out that, by choosing the particular complex Gaussian initial condition $\widehat{\mathbf{u}}^{(1),\nu}(0,\mathbf{k})$ to coincide with a fractional Gaussian vector $\widehat{\mathbf{u}}^{\text{\tiny{fGv}},\nu}(\mathbf{k}_{\mathbf{n}})$ as depicted in subsection 3.2.2, then, it the initial condition is fully characterized by the following covariances
$\mathbb{E}\left[\widehat{u}^{(1),\nu}_{i}(0,\mathbf{k}_{\mathbf{n}})\widehat{u}^{(1),\nu}_{j}(0,\mathbf{k}_{\mathbf{m}})\right]=\frac{L^{3}_{\text{\tiny{tot}}}}{2}E_{\nu}^{\text{\tiny{E}}}(k_{\mathbf{n}})\widehat{P}_{ij}(\mathbf{k}_{\mathbf{n}})\delta_{\mathbf{n},-\mathbf{m}}^{(3)},$ (76)
and
$\mathbb{E}\left[\widehat{u}^{(1),\nu}_{i}(0,\mathbf{k}_{\mathbf{n}})\overline{\widehat{u}^{(1),\nu}_{j}(0,\mathbf{k}_{\mathbf{m}})}\right]=\frac{L^{3}_{\text{\tiny{tot}}}}{2}E_{\nu}^{\text{\tiny{E}}}(k_{\mathbf{n}})\widehat{P}_{ij}(\mathbf{k}_{\mathbf{n}})\delta_{\mathbf{n},\mathbf{m}}^{(3)},$ (77)
which can be written as
$\widehat{u}^{(1),\nu}_{i}(0,\mathbf{k}_{\mathbf{n}})=\sqrt{\frac{E_{\nu}^{\text{\tiny{E}}}(k_{\mathbf{n}})}{T_{\mathbf{k}_{\mathbf{n}}}}}\widehat{P}_{ip}(\mathbf{k}_{\mathbf{n}})\int_{s=-\infty}^{0}e^{\frac{s}{T_{\mathbf{k}_{\mathbf{n}}}}}\widehat{{\mathrm{d}}W}_{p}(s,\mathbf{k}_{\mathbf{n}}),$ (78)
allows us to write the unique statistically stationary solution of (74) as
@@
Let us now explore the consequences of the particular temporal kernel $F^{(1)}$ given in (81) on the regularity in time of the corresponding velocity field, and more precisely on the scaling behavior of the respective temporal structure function $S_{2}^{(1),\text{\tiny{T}}}(\tau)$, i.e. the variance of the velocity time increment $\delta_{\tau}\mathbf{u}^{(1),\nu}$ (39). Recall that we found the expression (2.4), that reads, using the generalized forms of $E^{\text{\tiny{E}}}(|\mathbf{k}|)$ (55) and $T_{\mathbf{k}}$, (56)
@@
Similarly to the characteristic time scale $\tau_{\text{c}}(|\ell|)$ (61), the temporal structure function $S_{2}^{(1),\text{\tiny{T}}}(\tau)$ (3.2.3) undergoes a transition at small scales depending on the values of $H$ and $\beta$. We have, while rescaling the dummy integration variable $\mathbf{k}$ by the appropriate power of the time scale $\tau$ for $H<\beta$:
$S_{2}^{(1),\text{\tiny{T}}}(\tau)$ $=\frac{1}{\pi}(1+2H)(3+2H)D_{2}|\tau|^{H/\beta}\int_{\mathbf{k}\in\mathbb{R}^{3}}\left[1-e^{-D_{3}|\mathbf{k}|_{|\tau|^{\frac{1}{2\beta}}/L}^{2\beta}}\right]|\mathbf{k}|^{2}|\mathbf{k}|_{|\tau|^{2\beta}/L}^{-(2H+5)}{\mathrm{d}}\mathbf{k}$
$\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\tau|\to 0}}4(1+2H)(3+2H)D_{2}|D_{3}\tau|^{H/\beta}\int_{\rho=0}^{\infty}\left[1-e^{-\rho^{2\beta}}\right]\rho^{-(2H+1)}{\mathrm{d}}\rho\mbox{ for }H<\beta.$ (83)
Note that when $H=1/3$ and $\beta=1/2$, one retrieves (41) from the general formulation (3.2.3). For $H>\beta$, one has
$S_{2}^{(1),\text{\tiny{T}}}(\tau)\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\tau|\to 0}}4(1+2H)(3+2H)D_{2}|D_{3}\tau|\int_{\rho=0}^{\infty}\rho^{4}\sqrt{\rho^{2}+1/L^{2}}^{2\beta-(2H+5)}{\mathrm{d}}\rho\mbox{ for }H>\beta.$ (84)
The scaling behaviors derived in (3.2.3) and (84) pertain to the inertial range. We would also like to explore the behavior of the structure function in the dissipative range, first by taking the limit of vanishing scales $\tau\to 0$, and secondly the limit of vanishing viscosity $\nu\to 0$. When the viscosity $\nu$ is not zero, instead of (3.2.3), we have
@@
3.2.4 Generalization to a causal and differentiable-in-time framework
In the Section above, we have described the statistical properties in time of the Markovian model of Chaves et al. (2003) which can be seen as an Ornstein-Uhlenbeck type of evolution for Fourier modes (74). Fourier modes of the velocity field $\widehat{u}^{(1),\nu}_{i}(t,\mathbf{k})$ at a given wave number $\mathbf{k}$ are not differentiable in time because they solve a first-order evolution equation driven by a noise which is white-in-time. As a consequence, the temporal paths of $\widehat{u}^{(1),\nu}_{i}(t,\mathbf{k})$ share the same regularity as the Brownian motion and are only Hölder continuous. This non-differentiable nature of Fourier modes has a counterpart in the physical space: the corresponding velocity field $u^{(1),\nu}_{i}(t,\mathbf{x})$ is not differentiable in time. This is consistent with the behavior of the temporal second-order structure functions $S_{2,\nu}^{(1),\text{\tiny{T}}}(\tau)$, which behaves proportionally to the time scale $\tau$ (85), even for a finite viscosity $\nu>0$. If the velocity field were differentiable, we would expect instead a quadratic behavior, i.e. $S_{2,\nu}^{(1),\text{\tiny{T}}}(\tau)=O_{\nu}(\tau^{2})$ where $O_{\nu}(\tau^{2})$ stands for a term of order $\tau^{2}$ with a viscous dependent multiplicative factor. The necessity of proposing a dynamic capable of achieving this goal in a causal way is of tremendous importance. To do so, we will exploit an idea first proposed by Sawford (1991) and then generalized to an infinitely differentiable framework by Viggiano et al. (2020) that consists of replacing the white noise entering in the dynamics (74) by a continuous random force, which is itself of Ornstein-Uhlenbeck type and/or itself governed by a dynamics stirred by a continuous random force, which would ensure the finiteness of the first and/or subsequent time derivatives.
@@ 3.2.4 Generalization to a causal and differentiable-in-time framework
Let us now explore the consequences of the particular temporal kernel $F^{(N)}$ given in (96) (resp. $F^{(\infty)}$ given in (97)) on the regularity in time of the corresponding velocity field $u^{(N),\nu}_{i}(t,\mathbf{x})$ (resp. $\mathbf{u}^{(\infty),\nu}_{i}(t,\mathbf{x})$) obtained via its Fourier series (47). This will be illustrated with the scaling behavior of the temporal structure function, i.e. the variance of the velocity time increment (39). Using the generalized forms of $E^{\text{\tiny{E}}}(k)$ (55) and $T_{k}$, (56), the expression (2.4) reads
and when $N\to\infty$
@@
Similarly to the temporal structure function (3.2.3) associated with the exponential kernel (81), $S_{2}^{(N),\text{\tiny{temp}}}(\tau)$ (3.2.4) and $S_{2}^{(\infty),\text{\tiny{temp}}}(\tau)$ (3.2.4) eventually undergo a transition at small scales, but for a different range of values of $H$ and $\beta$. We indeed obtain,
Similarly for $S_{2}^{(\infty),\text{\tiny{temp}}}(\tau)$, replacing $F^{(N)}$ by $F^{(\infty)}$ in (3.2.4) and computing the remaining integral using the Gaussian correlation function (97), one obtains
Note that, for a given $\beta$, the range of accessible spatial regularity $H$ is wider (i.e. $H<2\beta$) in (3.2.4) and (102) than what was found (i.e. $H<\beta$) in (3.2.3). This is a consequence of the behavior of $1-F^{(N)}(x)$ and $1-F^{(\infty)}(x)$ near the origin which go to zero as $x^{2}$, while the exponential kernel (81) predicts a much slower decay (i.e. as $x$). This is shown in (146).
For a given $\beta$, when $H$ gets large, one obtains the trivial scaling associated to a differential field, using the expansion of the function $F^{(N)}$ provided in (146):
A similar result is satisfied for $S_{2}^{(\infty),\text{\tiny{temp}}}(\tau)$ as a consequence of (3.2.4) when $N\to\infty$.
The scaling behaviors derived in (3.2.4) and (3.2.4) pertain to the inertial range, for which the viscosity has been set to 0 before considering any limiting behavior. We now explore the behavior of the structure function taking first the limit of vanishing scales $\tau\to 0$, and second the limit of vanishing viscosity $\nu\to 0$. When the viscosity $\nu$ is not taken to zero, instead of (3.2.4), we have
In the case $H<2\beta$, we also observe a transition in the power-law exponent of the temporal structure function, taking value $S_{2}^{(N),\text{\tiny{temp}}}(\tau)\propto|\tau|^{H/\beta}$ in the inertial range (3.2.4), and then $S_{2,\nu}^{(N),\text{\tiny{temp}}}(\tau)\propto|\tau|^{2}$ in the dissipative range (3.2.4). This transition takes place at a dissipative time scale $\tau_{d}$ defined as the very characteristic time scale at which the asymptotic expansions of both $S_{2}^{(N),\text{\tiny{temp}}}$ (obtained from (3.2.4) ) and $S_{2,\nu}^{(N),\text{\tiny{temp}}}$ (obtained from (3.2.4)) coincide. Then it is clear that $\tau_{d}$ is fully determined by viscosity and can be shown to be related to a certain power (depending on $\beta$) of the dissipative length scale $\eta_{d}$ defined in the longitudinal PSD (54).
$\widehat{F}^{(\infty)}(\omega)=\sqrt{\pi}e^{-\pi^{2}\omega^{2}},$ (107)
We show in Fig. 4(b) the convergence of the 1D Fourier modes to the Gaussian correlation function. Details are explained in the Appendices A and C Letting first viscosity vanish ($\nu\to 0$), and considering the large frequencies $\omega\to\infty$ (secondly), we obtain for $H<(2N-1)\beta$
It is worth noting that, when $N$ is finite, for $H>(2N-1)\beta$, one has
Equations (3.2.4) and (109) show that, when $H<2\beta$, even if we let the number of layers $N$ tends to infinity, the field is not differentiable in time, as it is expected in the limit of vanishing viscosity. This range of possible values of the parameter $H$ setting the regularity in time is only accessible when $N\geq 2$. Depending on the values of $N$ and $\beta$, the present approach allows to consider Hurst parameters $H$ bigger that unity, that gives rise to fields at least once differentiable in time, although physically not really realistic in a turbulent context. When $H$ gets really large, and larger than $(2N-1)\beta$, the equivalent provided in (3.2.4) says that the field is differentiable (N1)(N-1)-times, the derivatives of order N1N-1 having the regularity of the Brownian motion.

### v2

@@ 3.2.1 The statistical structure of the stationary solution: spatial Fractional Gaussian vector Field
Let us first consider the purely spatial white noise ${\mathrm{d}}\mathbf{W}(\mathbf{x})$”, with (Hermitian symmetric) Fourier modes $\widehat{{\mathrm{d}}\mathbf{W}}(\mathbf{k}_{\mathbf{n}})$ covariance given by
$\mathbb{E}\left[\widehat{{\mathrm{d}}W}_{i}(\mathbf{k}_{\mathbf{n}})\widehat{{\mathrm{d}}W}_{j}(\mathbf{k}_{\mathbf{m}})\right]=L^{3}_{\text{\tiny{tot}}}\delta_{i,j}\delta_{\mathbf{n},-\mathbf{m}}^{(3)}.$ (62)
The real random field $\mathbf{u}^{\text{\tiny{fGv}},\nu}(\mathbf{x})$ is fully defined by its Fourier mode, and is given, component-wise, by
@@
for a given discrete wave vector $\mathbf{k}=\mathbf{k}_{\mathbf{n}}$. $T_{\mathbf{k}}$ is the characteristic time of the Fourier mode $\mathbf{k}$ defined in (53), $\widehat{P}_{ij}(\mathbf{k})$ Leray projector ensuring the incompressibility defined in (7) and $E_{\nu}^{\text{\tiny{E}}}(k)$ the previously defined 3D-PSD obtained from (51) using the isotropic relation (9). The superscript (1) appearing in the notation of the resulting velocity field $\widehat{u}^{(1),\nu}_{i}(t,\mathbf{k})$ (65) illustrates that it is obtained using a single evolution with respect to the white noise. The justification for this notation will become clear in the next paragraph. The random contribution to the evolution (65) is provided by the stochastic vector term $\widehat{{\mathrm{d}}W}_{p}(t,\mathbf{k})$ characterized by the covariance (61). Let us mention a very recent work by Letournel et al. (2025) devoted to an equivalent formulation of Chaves et al. (2003) for a decomposition of the velocity field in a wavelet dyadic basis.
The unique statistically stationary solution of the stochastic dynamics proposed in (65) reads
@@
Let us now explore the consequences of the particular temporal kernel $F^{(1)}$ given in (68) on the regularity in time of the corresponding velocity field, and more precisely on the scaling behavior of the respective temporal structure function $S_{2}^{(1),\text{\tiny{T}}}(\tau)$, i.e. the variance of the velocity time increment $\delta_{\tau}\mathbf{u}^{(1),\nu}$ (39).
Recall that we found the expression (2.4), that reads, using the generalized forms of $E^{\text{\tiny{E}}}(|\mathbf{k}|)$ (52) and $T_{\mathbf{k}}$, (53)
@@
Similarly to the characteristic time scale $\tau_{\text{c}}(\ell)$ (58), the temporal structure function $S_{2}^{(1),\text{\tiny{T}}}(\tau)$ (3.2.2) undergoes a transition at small scales depending on the values of $H$ and $\beta$. We have, while rescaling the dummy integration variable $\mathbf{k}$ by the appropriate power of the time scale $\tau$ for $H<\beta$:
$S_{2}^{(1),\text{\tiny{T}}}(\tau)$ $\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\tau|\to 0}}4(1+2H)(3+2H)D_{2}|D_{3}\tau|^{H/\beta}\int_{\rho=0}^{\infty}\left[1-e^{-\rho^{2\beta}}\right]\rho^{-(2H+1)}{\mathrm{d}}\rho.$ (70)
Note that when $H=1/3$ and $\beta=1/2$, one retrieves (41) from the general formulation (70). For $H>\beta$, one has
$S_{2}^{(1),\text{\tiny{T}}}(\tau)\mathrel{\mathop{\kern 0.0pt\sim}\limits_{|\tau|\to 0}}4(1+2H)(3+2H)D_{2}|D_{3}\tau|\int_{\rho=0}^{\infty}\rho^{4}\sqrt{\rho^{2}+1/L^{2}}^{2\beta-(2H+5)}{\mathrm{d}}\rho.$ (71)
The scaling behaviors derived in (70) and (71) pertain to the inertial range. We would also like to explore the behavior of the structure function in the dissipative range, first by taking the limit of vanishing scales $\tau\to 0$, and secondly the limit of vanishing viscosity $\nu\to 0$. When the viscosity $\nu$ is not zero, instead of (3.2.2), we have
@@
3.2.3 Generalization to a Markovian and differentiable-in-time framework
In the Section above, we have described the statistical properties in time of the Markovian model of Chaves et al. (2003) which can be seen as an Ornstein-Uhlenbeck type of evolution for Fourier modes (65). Fourier modes of the velocity field $\widehat{u}^{(1),\nu}_{i}(t,\mathbf{k})$ at a given wave number $\mathbf{k}$ are not differentiable in time because they solve a first-order evolution equation driven by a noise which is white-in-time. As a consequence, the temporal paths of $\widehat{u}^{(1),\nu}_{i}(t,\mathbf{k})$ share the same regularity as the Brownian motion and are only Hölder continuous. This non-differentiable nature of Fourier modes has a counterpart in the physical space: the corresponding velocity field $u^{(1),\nu}_{i}(t,\mathbf{x})$ is not differentiable in time. This is consistent with the behavior of the temporal second-order structure functions $S_{2,\nu}^{(1),\text{\tiny{T}}}(\tau)$, which behaves proportionally to the time scale $\tau$ (72), even for a finite viscosity $\nu>0$. If the velocity field were differentiable, we would expect instead a quadratic behavior, i.e. $S_{2,\nu}^{(1),\text{\tiny{T}}}(\tau)=O_{\nu}(\tau^{2})$ where $O_{\nu}(\tau^{2})$ stands for a term of order $\tau^{2}$ with a viscosity-dependent multiplicative factor. The necessity of proposing a dynamic capable of achieving this goal in a Markovian way is of tremendous importance, both from a physical and numerical perspectives. To do so, we will exploit an idea first proposed by Sawford (1991) and then generalized to an infinitely differentiable framework by Viggiano et al. (2020) that consists of replacing the white noise entering in the dynamics (65) by a continuous random force, which is itself of Ornstein-Uhlenbeck type and/or itself governed by a dynamics stirred by a continuous random force, which would ensure the finiteness of the first and/or subsequent time derivatives.
@@ 3.2.3 Generalization to a Markovian and differentiable-in-time framework
Let us now clarify the Markovian nature of the evolution proposed in the set of equations (77) to (80). Although the very evolution of the respective Fourier mode $\widehat{u}^{(N),\nu}_{i}(t,\mathbf{k})$ cannot be considered as Markovian in itself, the $N$-dimensional vector made up of that Fourier mode and the (N1)(N-1) remaining underlying layers $\widehat{f}^{(m)}_{i}(t,\mathbf{k})$ for $N-1\geq m\geq 1$, as it is defined in (101) considered in the numerical Section 4, can be formulated in a Markovian way. The respective matrix evolution is given in (104).
@@
Surprisingly, the expression of the correlation function $F^{(N)}$ (83) coincides exactly with the one of a Matérn Gaussian process, which have been extensively used in machine learning theory, as reviewed by Williams & Rasmussen (2006). More recently, such Gaussian processes have been exploited in a turbulent modeling context by Lilly et al. (2017). To our understanding, this type of covariance has been called in these formerly cited investigations for the necessity of considering differentiable Gaussian processes. Our approach furthermore shows that this covariance function of Matérn type (83) can be realized in a Markovian fashion, following the set of embedded Ornstein-Uhlenbeck evolutions provided in equations (77) to (80). In this sense, it can be viewed as a new way to interpret the nature of Matérn Gaussian processes.
@@
Let us now explore the consequences of the particular temporal kernel $F^{(N)}$ given in (83) (resp. $F^{(\infty)}$ given in (84)) on the regularity in time of the corresponding velocity field $u^{(N),\nu}_{i}(t,\mathbf{x})$ (resp. $u^{(\infty),\nu}_{i}(t,\mathbf{x})$}). We gather in Appendix (B) the detailed derivation necessary to justify them.
Let us first focus on the temporal structure functions, i.e. the variance of the velocity time increment (39). Using the generalized forms of $E^{\text{\tiny{E}}}(k)$ (52) and $T_{k}$, (53), the expression (2.4) reads
$S_{2}^{(N),\text{\tiny{temp}}}(\tau)$ $=\lim_{L_{\text{\tiny{tot}}}\to\infty}\lim_{\nu\to 0}\mathbb{E}\left[\left|\delta_{\tau}\mathbf{u}^{(N),\nu}\right|^{2}\right],$ (86)
which includes the limiting case $N=\infty$.
@@
Similarly to the temporal structure function (3.2.2) associated with the exponential kernel (68), $S_{2}^{(N),\text{\tiny{temp}}}(\tau)$ (86) eventually undergoes a transition at small scales, but for a different range of values of $H$ and $\beta$. We indeed obtain for any $N$, including $N=\infty$,
$S_{2}^{(N),\text{\tiny{temp}}}$ $(\tau)\mathrel{\mathop{\kern 0.0pt\propto}\limits_{|\tau|\to 0}}|D_{3}\tau|^{H/\beta}\mbox{ for }H<2\beta.$ (87)
We provide in (B) and (130) an explicit expression of the multiplicative constant. Note that, for a given $\beta$, the range of accessible spatial regularity $H$ is wider (i.e. $H<2\beta$) in (87) than what was found (i.e. $H<\beta$) in (70). The behavior of the structure function for $H>2\beta$ is provided in (B).
Importantly, seeking for the limiting behavior at small scales $\tau\to 0$ without taking first the vanishing viscosity limit, we now obtain as soon as $N\geq 2$
$S_{2,\nu}^{(N),\text{\tiny{temp}}}(\tau)\mathrel{\mathop{\kern 0.0pt\propto}\limits_{|\tau|\to 0}}\left|D_{3}\tau\right|^{2}.$ (88)
Once again, a more precise expression than (88) is given in (B). Thus, it now behaves proportionally to $\tau^{2}$ for any pair of values of $H$ and $\beta$: the obtained velocity field $\mathbf{u}^{(N),\nu}(t,\mathbf{x})$ is now smooth in time as long as $\nu>0$ and $N\geq 2$.
Considering now the associated temporal spectrum $E_{\nu}^{(N),\text{\tiny{T}}}(\omega)$ defined by
$E_{\nu}^{(N),\text{\tiny{T}}}(\omega)$ $=\lim_{L_{\text{\tiny{tot}}}\to\infty}\int_{\tau\in\mathbb{R}}e^{-2i\pi\omega\tau}\mathbb{E}\left[\mathbf{u}^{(N),\nu}(t,\mathbf{x})\cdot\mathbf{u}^{(N),\nu}(t+\tau,\mathbf{x})\right]{\mathrm{d}}\tau,$ (89)
we obtain for $H<(2N-1)\beta$
$E^{(N),\text{\tiny{T}}}(\omega)$ $\equiv\lim_{\nu\to 0}E_{\nu}^{(N),\text{\tiny{T}}}(\omega)\mathrel{\mathop{\kern 0.0pt\propto}\limits_{\omega\to\infty}}\omega^{-\left(1+H/\beta\right)},$ (90)
with a more precise expression given in (B), and in (137) for the limiting case $N\to\infty$, The case $H>(2N-1)\beta$ is treated in (B).